OpenStudy (anonymous):

Which proportion is true? A. 4/10=3/6 B. 1/2=7/8 C. 1/2=3/6 D. 4/10=7/8

OpenStudy (anonymous):

1/2 =3/6 1x6 = 3x2 6=6

OpenStudy (anonymous):

C 1/2 literally means half, just like 3 means half to 6.

OpenStudy (anonymous):

or (3:3/6:3) = 1/2 The answer is C

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Welcome^_^